OpenStudy (anonymous):

Write the equation of the line that passes through points C and D C(-1,4) and D(6,-6)

jimthompson5910 (jim_thompson5910):

First find the slope of the line through the two points. m = (y2-y1)/(x2-x1) m = (-6-4)/(6-(-1)) m = (-6-4)/(6+1) m = (-10)/7 m = -10/7 So the slope of the line that passes through the two points is -10/7 Now use y = mx+b along with (x,y) = (-1,4) to find the equation of the line. \[\Large y = mx+b\] \[\Large 4 = \left(\frac{-10}{7}\right)\left(-1\right)+b\] \[\Large 4 = \frac{10}{7}+b\] \[\Large 4-\frac{10}{7} = b\] \[\Large \frac{18}{7} = b\] \[\Large b = \frac{18}{7}\] So the equation of the line is \[\Large y = -\frac{10}{7}x+\frac{18}{7}\]